OpenStudy (anonymous):

1. Solve for x in the interval [0 , 2 π ). (a) 2sin^2x=sin x (b) 3sin^2x+cos^2x=2 (c) 1+3cosx=cos2x (d) sin2x+sin x=0

OpenStudy (anonymous):

a) rearrange to get 2sin^2x-sinx=0 then factor sinx(2sinx-1)=0 then set both sinx and 2sinx-1 equal to 0 and solve sinx=o means at what angle is the sin 0, this is 0 2sinx-1=0 rearrange to get sinx=1/2 at what angle is the sin 1/2, this is pi/6 and 5pi/6, I believe.

OpenStudy (anonymous):

b)using the identity sin^2x +cos^2x=1 rearranged to cos^2x=1-sin^2x 3sin^2x+1-sin^2x=2 combining like terms we get 2sin^2x=1 solving for sinx we get sinx=+-sqrt(2)/2 At what angle is the sin +-sqrt(2)/2? That would be pi/4, 3pi/4, 5pi/4 and 7pi/4

OpenStudy (anonymous):

c) cos2x=2cos^2 x - 1 substituting in we get 1+3cosx=2cos^2x-1 Rearranging we get 2cos^2x-3cosx-2=0 Factoring we get (2cosx+1)(cosx-2)=0 setting both = to 0 gives cosx=-1/2 and cosx=2 Now Im not sure what to do about the cosx=2 part but what angle has a cos of -1/2? That would be 2pi/3 and 4pi/3, I think.

OpenStudy (anonymous):

d) using the double angle formula sin2x=2sinxcosx we substitute and get 2sinxcosx+sinx=0 Factoring out a sinx we get sinx(2cosx+1)=0 setting both = to 0 we get sinx=0 and 2cosx+1=0 what angle in the interval has a sin of 0 and a cos of -1/2 this is 0, pi (I forgot this solution in part a), 2pi/3 and 4pi/3

OpenStudy (anonymous):

Hope this helps. Please check the math as I did it rather quickly and somewhat distractedly :)

OpenStudy (anonymous):

have cjecked it, perfect thanks :)